# Large-Scale Multi-dimensional Knowledge Profiling of Scientific Literature

### v1

@@ 1 Introduction
Existing tools only partially address this need. Traditional bibliometric methods, which built on metadata, co-citation networks, and keyword statistics [43, 16, 18], provide high-level overviews but capture limited semantic information and treat topics as largely static entities. As a result, they offer only coarse insight into methodological transitions, cross-domain influences, or the finer-grained structure of research problems. Meanwhile, recent systems that incorporate large language models (LLMs) [7, 1] demonstrate improved semantic analysis, supporting tasks such as retrieval-augmented question answering [17] and automated survey generation [37]. However, these tools are typically designed for short-range retrieval, single papers, or narrow tasks, and they do not provide a coherent, longitudinal view of large scientific corpora.
@@
Figure 2: LLM-Driven Multi-Dimensional Knowledge Profiling Framework with Large-Scale Literature Database and Hierarchical Retrieval.
These gaps point to the need for a unified way to organize, summarize, and interpret the rapidly expanding body of AI literature. In this work, we construct a large-scale profiling pipeline aimed at characterizing the recent development of AI research. Using more than 100,000 papers from 22 major conferences published between 2020 and 2025 (Figure 1), we combine text clustering, LLM-assisted semantic parsing, and lightweight retrieval techniques to form a structured representation of research problems, methods, datasets, and topical dynamics. Rather than emphasizing algorithmic novelty, our focus is on creating a coherent analytic framework that enables researchers to explore and reason about the field at multiple levels of granularity.
@@
2 Related Work
2.1 Bibliometrics & Scientific Trend Analysis
Content mining and trend analysis of scientific literature are important topics in scientometrics and information science. Traditional bibliometric approaches rely on metadata, extracting abstracts, authors, journals, keywords, and citations[43, 16], and applying co-citation and co-word analyses[35, 10, 3] to reveal the structure and evolution of research fields. However, these methods mainly rely on surface features and are limited in capturing the latent semantic structures within documents. To address this, topic modeling and unsupervised clustering have been widely used[6], with LDA inferring topic distributions from word co-occurrences[5, 2, 9, 12]. More recently, embedding-based clustering methods, such as BERTopic[18, 29] combined with BERT[14], and applications of large language models[20, 15, 8, 22] have enhanced the discovery of latent topics. Nevertheless, existing methods still struggle to systematically represent the dynamic evolution of knowledge[31, 30], and a unified framework for multidimensional knowledge profiling remains lacking.
2.2 LLMs for Scientific Document Understanding
Trained on massive general-domain corpora, LLMs have shown strong abilities in semantic understanding and text generation [7], giving rise to series of GPT [1] and other state-of-the-art models [39, 13, 34, 42]. To support scientific applications, researchers have explored LLMs in research workflows. For example, PaperQA [21, 33, 4] leverages retrieval-augmented generation (RAG) techniques [40, 17, 23, 11] to enable document question answering, while multi-agent and reinforcement learning approaches [28, 32, 23, 38] further improve system accuracy. Another line of work focuses on automating literature review generation, including AutoSurvey [37], SurveyX[24], SurveyForge [24, 41], and SurveyG [27], which implement pipelines for retrieval, filtering, organization, and writing. Despite these advances, existing LLM-based approaches are generally limited in scope: they focus on single tasks, rely on predefined corpora, and seldom provide a unified view of scientific knowledge evolution across multiple dimensions, including method evolution, task evolution, dataset adoption, and compute trends. Moreover, few approaches systematically integrate temporal information, semantic structure, and cross-conference knowledge, limiting their utility for large-scale trend analysis and evidence-based research decision making.
@@ 2.2 LLMs for Scientific Document Understanding
3 Knowledge Profiling Framework
We collected papers accepted between 2020 and 2025 from 22 academic conferences, including CVPR, ECCV, ICCV, ICLR, NeurIPS, and ICML, along with the citation counts for each paper, and over 100,000 papers are included in total. Our main focus is on how to efficiently perform Multi-Dimensional Knowledge Profiling on this scientific data, as illustrated in Fig. 2.
3.1 Semantic Parsing & Knowledge Extraction
To systematically analyze the rapidly growing body of AI research literature, we propose a multi-stage knowledge extraction framework that combines automated content mining with LLM-driven semantic analysis. First, considering the difficulties in structured output and efficiency when directly using LLMs to analyze PDF files, we employ minerU [36] to convert the collected PDFs into structured Markdown files. Subsequently, based on the Markdown file corresponding to each paper, we utilize Deepseek-R1-32B [19] to conduct multi-dimensional analysis of the papers, including meta information, abstract, research questions, methods, datasets, evaluation metrics, model architecture, contributions, limitations, and future directions.
It is worth noting that in multi-dimensional content parsing, the topic information of the papers is crucial. However, when LLMs process large scale structured data, on the one hand, due to the hallucination problem, LLMs struggle to stably output topic classifications, and on the other hand, the processing speed is limited. To address this, we adopt a clustering method similar to BERTopic[18] for topic clustering of papers. The specific process is as follows: first, use text encoder to process the title and abstract of each paper, then perform dimensionality reduction via UMAP[26], and then use HDBSCAN[25] for clustering, resulting in more than 300 topic categories distinguished based on semantic similarity. Finally, we use ChatGPT-5 to generate topic summaries for each cluster, extract core topic names, and construct a hierarchical logical relationship of topics, forming a multi-dimensional knowledge database, $ResearchDB$, which can be used for subsequent analysis. Unlike traditional bibliometric methods, this framework relies on semantic understanding rather than mere keywords, enabling the discovery of emerging or previously underexplored research topics, and finally obtaining the following information:
@@ 3.1 Semantic Parsing & Knowledge Extraction
This data and code will be open soon.
3.2 Intent-Driven Knowledge Retrieval
@@
Hierarchical Retrieval Each sub-question $Q_{i}$ is addressed through a two-level retrieval strategy:
@@
    Metadata Filtering. We filter $ResearchDB$ using metadata attributes such as conference, year, authors, and keywords to obtain a candidate subset:
@@
    Weighted Multi-Field Semantic Search. The content fields (abstract, methods, datasets, and limitations) of documents in $D_{\text{filtered}}$ are mapped to semantic vectors. Given $s_{ij}$ as the similarity score for field $j$ with sub-question $Q_{i}$, and $w_{j}$ as the field weight, the final relevance score for document $i$ is:
@@
This pipeline improves over direct language model generation by explicitly modeling user intent, combining coarse metadata filtering with weighted multi-field semantic search for scalable retrieval, and producing transparent, evidence-based answers with traceable provenance. It supports complex, multi-dimensional queries, facilitates analysis across topics, datasets, and methods, and scales effectively to large scientific corpora.
4 Scientific Insights and Findings
@@
4.1 Topic Growth and Transitions
@@
By analyzing the distribution of various topics in the four quadrants, as shown in Fig. 3, we observe that AI research is accelerating its evolution from being driven by early ”model scale” and ”basic perception tasks” to a paradigm driven by ”safety and controllability, multimodal cognition, and intelligent agent systems”. To further verify the overall trend of the life cycle analysis, we counted the changes in the number of papers on various topics over the past five years, and aslo selected the Top-20 topics with the largest number of papers for time series analysis, as shown in the lower left in Fig. 3.
@@
4.2 Compute & Model Scale Analysis
With the rise of large models and the increase in complex tasks, the consumption of computing resources and the scale of model parameters have become important dimensions for measuring research trends. We analyzed the information on the use of computing resources in the statistical data and unified all computing power into A100 Equivalent Hours, as shown in Fig. 4. The analysis results show that the consumption of training resources has maintained a significant upward trend, especially in related research such as generative models, multimodal models, and intelligent agents. This trend is consistent with the current pursuit of model scale and complexity in the AI field, reflecting the reliance of higher-performance models on huge computing power.
@@ 4.2 Compute & Model Scale Analysis
Figure 5: Dataset and Benchmark Trends (2020-2025). The bar chart shows annual paper counts for each dataset type, and the line chart shows their proportion of total publications.
@@
4.3 Dataset Landscape
Figure 6: Topic distribution in Dataset/Benchmark papers. The most frequent topics are labeled.
Benchmark papers lay a crucial foundation for the standardization and comparability of technological progress by providing unified evaluation criteria and datasets. Statistics from the sample set show that the proportion of Benchmark papers is on a significant upward trend, with the growth rate accelerating particularly after 2022, as shown in Fig. 5. Their proportion rose rapidly from 10.07% in 2022 to 18.16% in 2025 (partial data), approaching one-fifth of the total number of papers. This explosive growth is mainly due to the rise of general artificial intelligence technologies represented by LLMs, as well as the academic community’s emphasis on the standardized evaluation of non-performance indicators such as model fairness, ethics, and robustness.
Further in-depth analysis of the themes of Benchmark articles reveals the evaluation challenges faced by current research hotspots and cutting-edge technologies. As shown in the statistical chart of theme count and proportion (Fig. 6), research focuses are highly concentrated on LLMs, multimodality, and complex reasoning tasks. The top ten themes are almost entirely centered around the latest advancements and applications of LLMs, fully reflecting the evaluation-driven characteristics of contemporary scientific research.
Against the backdrop of this surge in benchmarks, the frequency of use of various datasets in scientific research papers has also shown rapid growth. This reflects the need for deeper evaluation in the era of LLMs and signifies that the paradigm for using datasets is no longer limited to single tasks but is evolving toward cross-domain and complex capability assessment, as shown in Fig. 7
@@ 4.3 Dataset Landscape
Figure 7: Trends in Representative Dataset Usage (2020-2025). The grouped bar chart shows annual paper counts per dataset, with overlaid lines highlighting trends for MMLU, GSM8K, ImageNet, COCO, and MATH.
4.4 Institution-Level Research Patterns
Figure 8: Publication Counts of Leading Institutions (2020-2025). Shows the number of papers from representative academic and industrial institutions across 22 conferences.
Figure 9: (a) Heatmap of topic-institution for top pairs; (b) Heatmap of institution-institution for top collaborations; (c) Word Clouds of Research Topics for Google; (d) Word Clouds of Research Topics for Tsinghua University.
Core institutions of China and the United States mainly maintained a dominant position across the period, as shown in Fig. 8. Tsinghua University has long been at the top in terms of the number of publications, demonstrating the stability of its research output and its ability to continuously accumulate. However, there were annual competitive dynamic changes in the rankings between industrial giants and universities. In terms of the cooperation network, as shown in Fig. 9, domestic universities have active collaborations with industrial giants such as Alibaba, Tencent, and Huawei, showing obvious regional characteristics. At the same time, Chinese and American institutions still dominate the cooperation network, forming a ”core-multipoint” cooperation pattern. This shift in research paradigms, driven by LLMs and Benchmark, is also profoundly reflected in the research strategies of different institutions. In the AI field, the academic community and the industrial sector have their own focuses and develop in a complementary manner. This is not simply a difference in interests, but rather reflects the strategic differentiation in the ecological niche of scientific research:
@@ 5 Retrieval and Analysis
The results in Fig. 10 indicate that the LLM-driven retrieval-generation method achieves higher accuracy than direct ChatGPT generation, particularly for topics spanning multiple sub-directions. By grounding its output in retrieved evidence from $ResearchDB$, the method minimizes hallucinations and ensures that the generated content is supported by verifiable sources. As a consequence, the richness of content may be somewhat lower compared to direct generation, and the text readability can be slightly reduced due to its more factual, evidence-focused style. Nonetheless, the approach excels in capturing emerging trends and providing reliable, evidence-based insights, making it a robust tool for large-scale literature analysis, research topic discovery, and informed decision-making in scientific exploration. Its ability to combine semantic retrieval with structured LLMs summarization offers a principled balance between precision and coverage, which is crucial for systematic, high-stakes analyses of evolving research domains.
Figure 10: Radar chart of ChatGPT-5 performance vs. ChatGPT-5 assisted by ResearchDBResearchDB, evaluated by human judges.
6 Conclusion
We present a multidimensional knowledge profiling framework combining topic clustering, LLM-based semantic parsing, and hierarchical retrieval to analyze over 100K AI papers. The framework captures topic dynamics, dataset and model trends, and institutional patterns, enabling evidence-driven discovery of emerging research directions, methodological shifts, and evolving technical paradigms. By integrating semantic understanding with structured retrieval, it provides both a macro-level overview of the research landscape and a fine-grained, topic-level perspective, supporting trend identification, research planning, and meta-analysis.
Limitations, broader impact and social impact. Our analysis relies on publicly accessible papers and metadata, which may be incomplete or biased, and LLM parsing can misinterpret domain-specific details. The study focuses on major English-language AI conferences, limiting generality. Despite these limitations, the framework provides a transparent, evidence-based view of research trends, though results should be interpreted cautiously to avoid overgeneralization or unintended influence.

### v2

# Multi-Dimensional Knowledge Profiling with Large-Scale Literature Database and Hierarchical Retrieval

@@ Introduction
Existing tools only partially address this need. Traditional bibliometric methods, which built on metadata, co-citation networks, and keyword statistics [zupic2015bibliometric, donthu2021conduct, grootendorst2022bertopic], provide high-level overviews but capture limited semantic information and treat topics as largely static entities. As a result, they offer only coarse insight into methodological transitions, cross-domain influences, or the finer-grained structure of research problems. Meanwhile, recent systems that incorporate large language models (LLMs) [brown2020language, achiam2023gpt] demonstrate improved semantic analysis, supporting tasks such as retrieval-augmented question answering [gao2023retrieval] and automated survey generation [wang2024autosurvey]. However, these tools are typically designed for short-range retrieval, single papers, or narrow tasks, and they do not provide a coherent, longitudinal view of large scientific corpora.
These gaps point to the need for a unified way to organize, summarize, and interpret the rapidly expanding body of AI literature. In this work, we construct a large-scale profiling pipeline aimed at characterizing the recent development of AI research. Using more than 100,000 papers from 22 major conferences published between 2020 and 2025 (Fig.˜1), we combine text clustering, LLM-assisted semantic parsing, and lightweight retrieval techniques to form a structured representation of research problems, methods, datasets, and topical dynamics.
@@
Rather than emphasizing algorithmic novelty, our focus is on creating a coherent analytic framework that enables researchers to explore and reason about the field at multiple levels of granularity.
@@
Related Work
Bibliometrics & Scientific Trend Analysis
Content mining and trend analysis of scientific literature are important topics in scientometrics and information science. Traditional bibliometric approaches rely on metadata, extracting abstracts, authors, journals, keywords, and citations[zupic2015bibliometric, donthu2021conduct], and applying co-citation and co-word analyses[van2010software, chen2006citespace, aria2017bibliometrix] to reveal the structure and evolution of research fields. However, these methods mainly rely on surface features and are limited in capturing the latent semantic structures within documents. To address this, topic modeling and unsupervised clustering have been widely used[blei2012probabilistic], with LDA inferring topic distributions from word co-occurrences[blei2003latent, alsumait2009topic, chang2009reading, chuang2013topic]. More recently, embedding-based clustering methods, such as BERTopic[grootendorst2022bertopic, reimers2019sentence] combined with BERT[devlin2019bert], and applications of large language models[kostikova2025lllms, diaz2025k, ccelikten2025topic, lam2024concept] have enhanced the discovery of latent topics. Nevertheless, existing methods still struggle to systematically represent the dynamic evolution of knowledge[si2024can, scherbakov2025emergence], and a unified framework for multidimensional knowledge profiling remains lacking.
LLMs for Document Understanding
Trained on massive general-domain corpora, LLMs have shown strong abilities in semantic understanding and text generation [brown2020language], giving rise to series of GPT [achiam2023gpt] and other state-of-the-art models [adakd, comanici2025gemini, touvron2023llama, yang2025qwen3]. To support scientific applications, researchers have explored LLMs in research workflows. For example, PaperQA [lala2023paperqa, skarlinski2024language, besrour2025squai] leverages retrieval-augmented generation (RAG) techniques [adavideorag, gao2023retrieval, li2025towards, cheng2025survey] to enable document question answering, while multi-agent and reinforcement learning approaches [nguyen2025ma, singh2025agentic, li2025towards, wu2025structure] further improve system accuracy. Another line of work focuses on automating literature review generation, including AutoSurvey [wang2024autosurvey], SurveyX[liang2025surveyx], SurveyForge [liang2025surveyx, yan2025surveyforge], and SurveyG [nguye2025surveyg], which implement pipelines for retrieval, filtering, organization, and writing. Despite these advances, existing LLM-based approaches are generally limited in scope: they focus on single tasks, rely on predefined corpora, and seldom provide a unified view of scientific knowledge evolution across multiple dimensions, including method evolution, task evolution, dataset adoption, and compute trends. Moreover, few approaches systematically integrate temporal information, semantic structure, and cross-conference knowledge, limiting their utility for large-scale trend analysis and evidence-based research decision making.
@@
Knowledge Profiling Framework
Figure 2: LLM-Driven Multi-Dimensional Knowledge Profiling Framework with Large-Scale Database and Hierarchical Retrieval.
Table 1: Detailed annual paper distribution for the 22 major venues in ResearchDBResearchDB from 2020 to 2025.
Venue Group 2020 2021 2022 2023 2024 2025 Total
Computer Vision (CVPR, ICCV, ECCV) 1358 3272 3715 4509 4868 2871 20593
Systems/Security (MLSYS, OSDI, NDSS, etc.) 216 374 468 467 479 36 2040
Overall Total (22 Venues) 7580 15324 16951 20944 25538 18145 104482
We collected papers accepted between 2020 and 2025 from 22 academic conferences, including CVPR, ECCV, ICCV, ICLR, NeurIPS, and ICML, along with the citation counts for each paper, and over 100,000 papers are included in total, as shown in Tab.˜1. Our main focus is on how to efficiently perform Multi-Dimensional Knowledge Profiling on this scientific data, as illustrated in Fig.˜2.
Semantic Parsing & Knowledge Extraction
To systematically analyze the rapidly growing body of AI research literature, we propose a multi-stage knowledge extraction framework that combines automated content mining with LLM-driven semantic analysis. First, considering the difficulties in structured output and efficiency when directly using LLMs to analyze PDF files, we employ minerU [wang2024mineru] to convert the collected PDFs into structured Markdown files. Subsequently, based on the Markdown file corresponding to each paper, we utilize Deepseek-R1-32B [guo2025deepseek] to conduct multi-dimensional analysis of the papers, including meta information, abstract, research questions, methods, datasets, evaluation metrics, model architecture, contributions, limitations, and future directions.
It is worth noting that in multi-dimensional content parsing, the topic information of the papers is crucial. However, when LLMs process large scale structured data, on the one hand, due to the hallucination problem, LLMs struggle to stably output topic classifications, and on the other hand, the processing speed is limited. To address this, we adopt a clustering method similar to BERTopic[grootendorst2022bertopic] for topic clustering of papers. The specific process is as follows: first, use text encoder to process the title and abstract of each paper, then perform dimensionality reduction via UMAP[mcinnes2018umap], and then use HDBSCAN[mcinnes2017hdbscan] for clustering, resulting in more than 300 topic categories distinguished based on semantic similarity. Finally, we use ChatGPT-5 to generate topic summaries for each cluster, extract core topic names, and construct a hierarchical logical relationship of topics, forming a multi-dimensional knowledge database, $ResearchDB$, which can be used for subsequent analysis. The ResearchDB schema captures five main dimensions (Summary, Technical, Analysis, System, Post-Processed) extracted from full paper texts using LLM-assisted parsing. Tab.˜2 provides the detailed documentation for the key fields in our structured data records. Unlike traditional bibliometric methods, this framework relies on semantic understanding rather than mere keywords, enabling the discovery of emerging or previously underexplored research topics, and finally obtaining the following information:
Table 2: Revised documentation for the key structural fields in the ResearchDB schema, reflecting the specific output of the LLM-assisted parsing pipeline.
Field Name Data Type Source Purpose/Description
conference/year String/Int Metadata Publication venue and year.
@@
Intent-Driven Knowledge Retrieval
Figure 3: The framework of Hierarchical Retrieval.
@@
Hierarchical Retrieval Each sub-question $Q_{i}$ is addressed through a two-level retrieval strategy, as shown inFig.˜3.
@@
    Metadata Filtering. We filter $ResearchDB$ using metadata attributes such as conference, year, authors, and keywords to obtain a candidate subset. Among them, keywords use fuzzy matching, and other fields use exact matching.
@@
    Weighted Multi-Field Semantic Search.
    For the document subset $D_{\text{filtered}}$ obtained via metadata filtering, we perform a subsequent intent-driven semantic search. Unlike traditional methods with static weights, our framework employs an adaptive weighting mechanism where the field weights $w_{j}$ are dynamically assigned by the LLM-based intent recognition module according to the query’s emphasis. For instance, if a sub-question $Q_{i}$ focuses on specific technical implementations, the module allocates significantly higher priority to the ’methods’ field (e.g., $w_{\text{methods}}=0.8$) relative to the ’abstract’ (e.g., $w_{\text{abstract}}=0.2$). These weights are then applied to the semantic similarity scores $s_{ij}$ calculated between the query and each core structural field—including abstract, methods, datasets, and limitations. The final weighted relevance score $S_{i}$ is computed as:
@@
Scientific Insights and Findings
@@
Topic Growth and Transitions
@@
By analyzing the distribution of various topics in the four quadrants, as shown in Fig.˜4, we observe that AI research is accelerating its evolution from being driven by early "model scale" and "basic perception tasks" to a paradigm driven by "safety and controllability, multimodal cognition, and intelligent agent systems". To further verify the overall trend of the life cycle analysis, we counted the changes in the number of papers on various topics over the past five years, and aslo selected the Top-20 topics with the largest number of papers for time series analysis, as shown in the lower left in Fig.˜4.
@@
Compute & Model Scale Analysis
With the rise of large models and the increase in complex tasks, the consumption of computing resources and the scale of model parameters have become important dimensions for measuring research trends. We analyzed the information on the use of computing resources in the statistical data and unified all computing power into A100 Equivalent Hours, as shown in Fig.˜5. The analysis results show that the consumption of training resources has maintained a significant upward trend, especially in related research such as generative models, multimodal models, and intelligent agents. This trend is consistent with the current pursuit of model scale and complexity in the AI field, reflecting the reliance of higher-performance models on huge computing power.
@@
(a) Dataset and Benchmark Trends (2020-2025). The bar chart shows annual paper counts for each dataset type, and the line chart shows their proportion of total publications.
(b) Topic distribution in Dataset/Benchmark papers. The most frequent topics are labeled.
Figure 6: The Analysis from the Dataset Landscape.
Dataset Landscape
Benchmark papers lay a crucial foundation for the standardization and comparability of technological progress by providing unified evaluation criteria and datasets. Statistics from the sample set show that the proportion of Benchmark papers is on a significant upward trend, with the growth rate accelerating particularly after 2022, as shown in Fig.˜6(a). Their proportion rose rapidly from 10.07% in 2022 to 18.16% in 2025 (partial data), approaching one-fifth of the total number of papers. This explosive growth is mainly due to the rise of general artificial intelligence technologies represented by LLMs, as well as the academic community’s emphasis on the standardized evaluation of non-performance indicators such as model fairness, ethics, and robustness.
Further in-depth analysis of the themes of Benchmark articles reveals the evaluation challenges faced by current research hotspots and cutting-edge technologies. As shown in the statistical chart of theme count and proportion (Fig.˜6(b)), research focuses are highly concentrated on LLMs, multimodality, and complex reasoning tasks. The top ten themes are almost entirely centered around the latest advancements and applications of LLMs, fully reflecting the evaluation-driven characteristics of contemporary scientific research.
Against the backdrop of this surge in benchmarks, the frequency of use of various datasets in scientific research papers has also shown rapid growth. This reflects the need for deeper evaluation in the era of LLMs and signifies that the paradigm for using datasets is no longer limited to single tasks but is evolving toward cross-domain and complex capability assessment, as shown in Fig.˜7(a)
@@
(a) Trends in Representative Dataset Usage (2020-2025). The grouped bar chart shows annual paper counts per dataset, with overlaid lines highlighting trends for MMLU, GSM8K, ImageNet, COCO, and MATH.
(b) Publication Counts of Leading Institutions (2020-2025). Shows the number of papers from representative academic and industrial institutions across 22 conferences.
Figure 7: Trends in Publication Counts for Dataset Usage and Leading Institutions from 2020 to 2025.
Institution-Level Research Patterns
Figure 8: (a)–(b) Top topic-institution and institution-institution heatmaps; (c)–(d) Research topic word clouds for Google and Tsinghua University.
Core institutions of China and the United States mainly maintained a dominant position across the period, as shown in Fig.˜7(b). Tsinghua University has long been at the top in terms of the number of publications, demonstrating the stability of its research output and its ability to continuously accumulate. However, there were annual competitive dynamic changes in the rankings between industrial giants and universities. In terms of the cooperation network, as shown in Fig.˜8, domestic universities have active collaborations with industrial giants such as Alibaba, Tencent, and Huawei, showing obvious regional characteristics. At the same time, Chinese and American institutions still dominate the cooperation network, forming a "core-multipoint" cooperation pattern. This shift in research paradigms, driven by LLMs and Benchmark, is also profoundly reflected in the research strategies of different institutions. In the AI field, the academic community and the industrial sector have their own focuses and develop in a complementary manner. This is not simply a difference in interests, but rather reflects the strategic differentiation in the ecological niche of scientific research:
@@
Figure 9: Performance comparison of ResearchDBResearchDB against the SurveyX method and ChatGPT-5 across five evaluation dimensions, based on expert scoring.
@@
The evaluation results in Fig.˜9 reveal that our proposed framework outperforms the baseline methods across several critical dimensions. Compared to direct generation by ChatGPT-5, our approach achieves significantly higher Accuracy and Usefulness. This improvement is primarily attributed to our intent-driven hierarchical retrieval, which grounds the model’s responses in structured evidence extracted from specific content fields. While SurveyX[liang2025surveyx] focuses on general automation, our framework excels in capturing technical specifics through weighted multi-field search. This evidence-grounded approach prioritizes high-fidelity sections (e.g., methods, datasets), ensuring superior reliability and traceability over unconstrained LLM generation , which is essential for high-stakes research analysis
Ablation Study
Table 3: Ablation results on retrieval performance based on expert scores.
Variant Acc. Cov. Nov. Read. Use.
w/o MF 4.12 3.85 4.00 4.10 3.90
w/o DW 3.86 3.82 3.70 4.05 3.75
SF 3.70 3.10 3.20 4.20 3.25
ResearchDB 4.50 3.90 4.20 4.00 4.40
To evaluate the contribution of each core component in our hierarchical retrieval pipeline, we conduct an ablation study by comparing the full ResearchDB framework against three degraded variants consisting of a configuration without metadata filtering (w/o MF), a version without dynamic weighting (w/o DW) using a uniform weight distribution, and a single-field baseline (SF) that restricts semantic search to abstracts only. As demonstrated by the experimental results in Tab.˜3, the full framework consistently achieves the highest performance across all metrics, with the most significant drop in accuracy occurring when the intent-driven dynamic weighting mechanism is disabled. Specifically, the removal of dynamic weighting prevents the system from effectively prioritizing task-relevant fields; Furthermore, the exclusion of metadata filtering leads to a noticeable decrease in precision as the system fails to exclude papers from irrelevant venues or timeframes that share similar semantic embeddings. Finally, the inferior performance of the single-field baseline confirms that relying solely on abstracts is insufficient for deep knowledge profiling, as essential information regarding datasets and limitations is often omitted or overly generalized in paper summaries, proving that the integration of structured multidimensional fields with adaptive weights is vital for high-fidelity literature analysis.
Conclusion
We present a multidimensional knowledge profiling framework that integrates topic clustering, LLM-based semantic parsing, and hierarchical retrieval to analyze over 100K AI publications. By capturing topic dynamics, dataset trends, and institutional patterns, the framework enables evidence-driven discovery of methodological shifts and technical paradigms. Our approach provides both a macro-level landscape overview and fine-grained, topic-level insights, effectively supporting systematic trend analysis and research planning.
Limitations and Impact. While our study is limited to major English-language AI conferences and subject to potential LLM parsing biases, it offers a transparent, evidence-based perspective on research evolution. Future work will aim to improve domain-specific parsing accuracy and extend the framework to a broader range of scientific venues.